# Emergent multi-loop nested point gap in a non-Hermitian quasiperiodic lattice

### v2

@@ II Model
[…]
@@ III Multi-loop nested point gaps
Figure 2: (a) The phase diagram of fractal dimension ${\Gamma}$ in the $\lambda-E$ plane. Eigenenergies in the complex plane with $\lambda=0.2$ (b), $\lambda=0.7$ (c), $\lambda=1$ (d) and $\lambda=1.7$ (e). (f) Arg[det($H-E_{b}$)] versus $\theta$, where we set $E_{b}=1,~{}1,~{}1,~{}-0.95$ [marked as rhombus in (b)-(e)]. Throughout, we set $L=610$, $q=0.2$$\alpha=377/610$ and $N=2$.
@@ III.1 Traditional point gap with $\omega=1$
Under such circumstance, the model is an integrable system. The critical point of Anderson transition is located at $\lambda_{c}=J$. The corresponding Lyapunov exponent $\gamma=\ln(\lambda/J)$ SLonghi2019 ; SLonghi2019a ; YLiu2021 . When $\lambda<J$ ($\lambda>J$), the system is extended (localized) and the eigenenergy spectrum $E(k)=2\cos(k)$ [$E(k)=\lambda e^{ik}+\lambda e^{-ik}$] with $k\in(0,2\pi]$, which corresponds to the trivial (topological) phase. In the topological phase, the maximum winding number $\omega=1$, which equals to the summation number of the GSM potential.
@@ III.2 The point gap with $\omega=2$
In Fig. 2(a) we exhibit the fractal dimensions $\Gamma$ versus $\lambda$ and $E$. By contrast to the case of traditional point gap, the multi-loop nested point gaps emerge here accompanied by MEs. In concrete terms, for the former, there is no ME, and only a clean critical point ($\lambda=J$) exists between the localized and extended states. For the latter, however, the potential of one more summation term cause the ME to be in the interval $1.20.48<\lambda<$ 1.2 [see Fig. 2(a)]. Furthermore, the localized states gradually invade the extended region until all eigenstates become localized with an increasing $\lambda$. Meanwhile, as the eigenspectrum changes from pure real to complex, the system’s PT symmetry is gradually broken [see Fig. 2(b)-(e)]. One can also see that the localized state forms two separated circles at the maximum and minimum eigenvalues, which are connected by the eigenvalues of the extended state. And when all the eigenstates are localized, the two circles fuse to form a multi-loop nested structure [see Fig. 2(e)]. Fig. 2(f) exhibits the argument of the determinants $H-E_{b}$ as a function of the flux $\theta$, where the corresponding base points $E_{b}$ are marked as a rhombus in Fig. 2(b)-(e). One can find that the Arg[det($H-E_{b}$)] is always zero with $\theta$ for the extended phase ($\lambda=0.2$), corresponding to the trivial phase with winding number $\omega=0$. Under the condition of $\lambda=0.7$ and $\lambda=1$, the Arg[det($H-E_{b}$)] forms a cycle with the $\theta$, corresponding to the topological phase with winding numbers $\omega=1$. Under the condition of complete localization ($\lambda=1.7$), a topological phase with high winding number $\omega=2$ emerges with an increasing $\theta$.
@@ III.3 The point gap with $\omega=3$
Combining the above results of $N=2$ and $N=3$, one can find that the winding number, the circle number of point gaps and the number of MEs are all dominated by the potential energy structure, i.e. the number of summation terms $N$. In concrete terms, the number of summation terms $N$ corresponds to the winding number $\omega=N$, the circle number of point gaps is equal to $N$, and the number of MEs is equal to $2(N-1)$.
@@ IV Theoretical analysis in infinite limit ($N={\infty}$)
[…]
@@
From the Fig. 4(a), as the quasiperiodic strength $\lambda$ increases, the first phase transition occurs at $E=2$, and the last phase transition occurs at $E=-2$ ($\lambda_{c}\approx 3.5$). After that, the system enters the localized phase completely. For the case of $-2<E<2$, the critical points are in the interval $3.50<\lambda_{c}<$ 3.5.
@@ V The point gap in dual space
Figure 5: (a) The dual space phase diagram. Eigenenergies in the complex plane with $\lambda=1.7$ under PBCs (b) and OBCs (c). (d) The dual space wave function of the eigenstate corresponding to an eigenenergy of $-0.9217$. The other parameters $L=610$, $\alpha=377/610$, $q=0.2$ and $N=2$.
@@ Appendix A Details for GSM lattice model
Note that, when $N\rightarrow\infty$, the term $\lambda{e}^{-{N}q}{e}^{({N}+1)i(2{\pi}{\alpha}{j}+{\theta})}$ tends to 00. then, one can obtain
@@ Appendix B The evolution of point gap with respect to $\lambda$
Appendix C Discussions on phase diagram of different $N$
In fact, while the number of MEs increases with an increasing $N$, the interval between MEs decreases versus $N$. Then, for the case of $N\rightarrow\infty$, the interval between two nearest-neighboring MEs approaches 00, and finally merges into one ME [see Fig. 8].
Figure 8: $\Gamma$ phase diagram on $\lambda-E$ plane with different $N$. Parameters $\lambda=1.9$, $L=610$, $\alpha=377/610$ and $q=0.2$.
The relationship between the ME and the parameter $N$ can be summarized as follows. Under the condition of $N=1$, the system exhibits a traditional topological point gap, in which there is no ME and only a standard critical point, namely $\lambda_{c}=J$. With the increase of parameter $N$, multiple MEs gradually appear in the system, and the number of MEs is $2(N-1)$, which is consistent with the previous prediction. Finally, as $N$ increases, the exponential decay term at large $N$ makes the MEs all merge together, while only one ME can be observed on the phase diagram. One can obtain the same conclusion from the point of view of topological point gap analysis, i.e., although the increase of point gaps’ circle promotes a higher winding number with an increasing $N$, the maximum winding number decreases to $1$ due to the merging of point gaps’ circles.
@@ Appendix F Experimental realization and measurement
where $\sigma^{\pm}=\frac{1}{2}(\sigma_{x}\pm i\sigma_{y})$ with Pauli matrices $\sigma_{x}$ and $\sigma_{y}$. The Hamiltonian (24) can be transformed to the GSM lattice by defining operator $c^{\dagger}_{j}=\sigma^{+}_{j_{x}}=\left|\uparrow\right>_{j_{x}}\left<\downarrow\right|_{j_{x}}$ at each site $j_{x}$, where |$=70P_{1/2}$ and |$=70S_{1/2}$ for 87Rb atoms [see Fig. 11]. The dipole-dipole interaction between Rydberg atoms is given by $J=\frac{d^{2}}{R^{3}}$, where $d$ represents the transition dipole moment between the two Rydberg levels, $R$ is the distance between two nearest neighboring sites, $70$ is the principal quantum number, and the energy unit $J=2\pi\times 7.53$MHz. Note that, since the non-nearest neighboring effect decays with distance $R^{\nu}$ ($\nu\geq 3$), one can safely ignore the corresponding terms.

### v3

@@ II Model
$\omega(E_{b})=\lim_{L\rightarrow\infty}\frac{1}{2\pi i}\int_{0}^{2\pi}d\theta\frac{\partial}{\partial\theta}\ln\left\{\mathrm{det}[H(\theta\rightarrow\frac{\theta}{L})-E_{b}]\right\},$ (4)
@@ III Multi-loop nested point gaps
[…]
@@ III.1 Traditional point gap with $\omega=1$
Under such circumstance, the model is an integrable system. The critical point of Anderson transition is located at $\lambda_{c}=J$. The corresponding Lyapunov exponent $\gamma=\ln(\lambda/J)$ SLonghi2019 ; SLonghi2019a ; YLiu2021 . When $\lambda<J$ ($\lambda>J$), the system is extended (localized) and the eigenenergy spectrum $E(k)=2\cos(k)$ $E(k)=\lambda e^{ik}+\lambda e^{-ik}$] with $k\in(0,2\pi]$, which corresponds to the trivial (topological) phase. In the topological phase, the maximum winding number $\omega=1$, which equals to the summation number of the GSM potential.
@@ III.2 The point gap with $\omega=2$
In Fig. 2(a) we exhibit the fractal dimensions $\Gamma$ versus $\lambda$ and $E$. By contrast to the case of traditional point gap, the multi-loop nested point gaps emerge here accompanied by MEs. In concrete terms, for the former, there is no ME, and only a clean critical point ($\lambda=J$) exists between the localized and extended states. For the latter, however, the potential of one more summation term cause the ME to be in the interval $0.48<\lambda<1.2$ [see Fig. 2(a)]. Furthermore, the localized states gradually invade the extended region until all eigenstates become localized with an increasing $\lambda$. Meanwhile, as the eigenspectrum changes from pure real to complex, the system’s PT symmetry is gradually broken [see Fig. 2(b)-(e)]. One can also see that the localized state forms two separated circles at the maximum and minimum eigenvalues, which are connected by the eigenvalues of the extended state. And when all the eigenstates are localized, the two circles fuse to form a multi-loop nested structure [see Fig. 2(e)]. Fig. 2(f) exhibits the argument of the determinants $H-E_{b}$ as a function of the flux $\theta$, where the corresponding base points $E_{b}$ are marked as a rhombus in Fig. 2(b)-(e). One can find that the Arg[det($H-E_{b}$)] is always zero with $\theta$ for the extended phase ($\lambda=0.2$), corresponding to the trivial phase with winding number $\omega=0$. Under the condition of $\lambda=0.7$ and $\lambda=1$, the Arg[det($H-E_{b}$)] forms a cycle with the $\theta$, corresponding to the topological phase with winding numbers $\omega=1$. Under the condition of complete localization ($\lambda=1.7$), a topological phase with high winding number $\omega=2$ emerges with an increasing $\theta$.
@@ III.3 The point gap with $\omega=3$
Combining the above results of $N=2$ and $N=3$, one can find that the winding number, the circle number of point gaps and the number of MEs are all dominated by the potential energy structure, i.e. the number of summation terms $N$. In concrete terms, the number of summation terms $N$ corresponds to the winding number $\omega=N$, the circle number of point gaps is equal to $N$, and the number of MEs is equal to 2(N1)2(N-1).
@@ IV Theoretical analysis in infinite limit ($N={\infty}$)
where $\left\|\cdot\right\|$ denotes the matrix norm. The LE can be divided into two parts, i.e., $\gamma(E)=\gamma_{A}(E)+\gamma_{B}(E)$, in which
@@
From the Fig. 4(a), as the quasiperiodic strength $\lambda$ increases, the first phase transition occurs at $E=2$, and the last phase transition occurs at E=2E=-2 ($\lambda_{c}\approx 3.5$). After that, the system enters the localized phase completely. For the case of 2<E<2-2<E<2, the critical points are in the interval $0<\lambda_{c}<3.5$.
@@ V The point gap in dual space
Figure 5: (a) The dual space phase diagram. Eigenenergies in the complex plane with $\lambda=1.7$ under PBCs (b) and OBCs (c). (d) The dual space wave function of the eigenstate corresponding to an eigenenergy of 0.9217-0.9217. The other parameters $L=610$, $\alpha=377/610$, $q=0.2$ and $N=2$.
@@ Appendix A Details for GSM lattice model
Note that, when $N\rightarrow\infty$, the term $\lambda{e}^{-{N}q}{e}^{({N}+1)i(2{\pi}{\alpha}{j}+{\theta})}$ tends to 0. then, one can obtain
@@ Appendix B The evolution of point gap with respect to $\lambda$
Appendix C Discussions on phase diagram of different NN
In fact, while the number of MEs increases with an increasing $N$, the interval between MEs decreases versus $N$. Then, for the case of $N\rightarrow\infty$, the interval between two nearest-neighboring MEs approaches 0, and finally merges into one ME [see Fig. 8].
Figure 8: $\Gamma$ phase diagram on $\lambda-E$ plane with different NN. Parameters $\lambda=1.9$, $L=610$, $\alpha=377/610$ and $q=0.2$.
The relationship between the ME and the parameter $N$ can be summarized as follows. Under the condition of $N=1$, the system exhibits a traditional topological point gap, in which there is no ME and only a standard critical point, namely $\lambda_{c}=J$. With the increase of parameter $N$, multiple MEs gradually appear in the system, and the number of MEs is 2(N1)2(N-1), which is consistent with the previous prediction. Finally, as $N$ increases, the exponential decay term at large $N$ makes the MEs all merge together, while only one ME can be observed on the phase diagram. One can obtain the same conclusion from the point of view of topological point gap analysis, i.e., although the increase of point gaps’ circle promotes a higher winding number with an increasing $N$, the maximum winding number decreases to $1$ due to the merging of point gaps’ circles.
@@ Appendix F Experimental realization and measurement
where $\sigma^{\pm}=\frac{1}{2}(\sigma_{x}\pm i\sigma_{y})$ with Pauli matrices $\sigma_{x}$ and $\sigma_{y}$. The Hamiltonian (24) can be transformed to the GSM lattice by defining operator $c^{\dagger}_{j}=\sigma^{+}_{j_{x}}=\left|\uparrow\right>_{j_{x}}\left<\downarrow\right|_{j_{x}}$ at each site $j_{x}$, where $\left|\uparrow\right>=70P_{1/2}$ and $\left|\downarrow\right>=70S_{1/2}$ for 87Rb atoms [see Fig. 11]. The dipole-dipole interaction between Rydberg atoms is given by $J=\frac{d^{2}}{R^{3}}$, where $d$ represents the transition dipole moment between the two Rydberg levels, $R$ is the distance between two nearest neighboring sites, $70$ is the principal quantum number, and the energy unit $J=2\pi\times 7.53$MHz. Note that, since the non-nearest neighboring effect decays with distance $R^{\nu}$ ($\nu\geq 3$), one can safely ignore the corresponding terms.